# The Oracle Complexity of Simplex-based Matrix Games: Linear Separability and Nash Equilibria

### v3

@@ 4.2 Oracle complexity with a two-sided oracle
5 Nash equilibria: simplex/simplex games
We now turn to consider matrix games in which the domain $\mathcal{W}\subset\mathbb{R}^{d}$ is the probability simplex:
$\max_{\mathbf{w}\in\Delta^{d-1}}~\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~.$ (3)
As previously discussed, the objective above corresponds to finding a Nash equilibrium in zero-sum bilinear games. In this section we will assume that each entry of $A$ is in [1,+1][-1,+1], which is the standard normalization for this setting (cf. Nemirovski, 2004; Hadiji et al., 2024).
Following previous works on algorithms for this problem, we consider lower bounds for the oracle complexity of Eq. (3), using a two-sided oracle. We note that for this problem, a two-sided oracle is exactly equivalent to a first-order (or gradient) oracle for Eq. (3), which returns the gradient of $\mathbf{p}^{\top}A\mathbf{w}$ with respect to some given $\mathbf{p},\mathbf{w}$ (namely, $A\mathbf{w}$ and $\mathbf{p}^{\top}A$). Thus, our lower bound automatically applies to solving Eq. (3) using a first-order oracle.
We prove the lower bound in two stages: First, we show by reduction that it is sufficient to prove an oracle complexity lower bounds for the $\ell_{1}$/simplex game
$\max_{\mathbf{w}\in\mathbb{R}^{d}:\|\mathbf{w}\|_{1}\leq 1}~\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~,$ (4)
namely where the simplex domain is enlarged to the entire $\ell_{1}$ ball. In the second stage, we prove such a lower bound for the $\ell_{1}$/simplex game (which of course, may be of independent interest).
Following this plan, we start by proving that an algorithm for solving (3) (over the simplex) can be readily converted to an algorithm for solving (4) (over the $\ell_{1}$ ball), with similar guarantees. Thus, to prove an oracle complexity lower bound for simplex/simplex games (3), it is sufficient to prove a lower bound for solving $\ell_{1}$/simplex games (4).
@@ 5 Nash equilibria: simplex/simplex games
The formal proof appears in the appendix. In a nutshell, the idea is that given a matrix $A\in[-1,+1]^{n\times d}$ and an algorithm for a simplex/simplex matrix game, we can feed the algorithm with the matrix $(A;-A)\in\mathbb{R}^{n\times 2d}$, and convert the algorithm’s output $\mathbf{w}$ (a vector in the (2d1)(2d-1)-simplex) to a $d$-dimensional vector $\mathbf{w}^{\prime}$ in the $\ell_{1}$ unit ball, so that $A\mathbf{w}^{\prime}=(A;-A)\mathbf{w}$, leading to similar guarantees. We note that the reduction does require us to modify the matrix width $d$ by a factor of $2$, but this will not affect our lower bound by more than a small constant factor (as the bound we will show applies to any sufficiently large $d$). We also note that the computational complexity of the two algorithms in the reduction are essentially identical.
We now turn to consider an oracle complexity lower bound for $\ell_{1}$/simplex games (4), using a two-sided oracle. Our main result is the following:
Theorem 5.2.
The following holds for some large enough universal constant $c>0$: For any $T>c$, suppose $n,d$ are sufficiently large so that $d>cT$ and $n>cT^{2}\log(Tdn)$. Then for any deterministic algorithm for solving Eq. (4), there exists a matrix $A\in[-1,+1]^{n\times d}$ satisfying
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{c\log(dn)}\cdot\frac{1}{T^{3/2}}~,$
In particular, by fixing some small enough $\epsilon>0$ and choosing $T$ such that the lower bound in the theorem is at least $\epsilon$, we can restate the bound as follows: If the number of iterations $T$ is less than $\Omega(\log(dn)^{-2/3}\epsilon^{-2/3})$, there exists a matrix $A$ such that the optimal value is at least $\epsilon$, yet the algorithm’s output has value less than 0, hence is suboptimal by at least $\epsilon$. As discussed previously, by Proposition 5.1, this lower bound for $\ell_{1}$/simplex automatically extends to simplex/simplex games, up to a small numerical constant which is absorbed in the $\Omega(\cdot)$ notation.
To prove the theorem, we utilize a construction very similar to the $\ell_{2}$ case in the previous section, except that the vectors need to be scaled differently, to satisfy the different constraints on $A$ and $\mathbf{w}$. Moreover, since now the constraint on $A$ is that each individual entry is in [1,+1][-1,+1], we need to choose each $\mathbf{u}_{t}$ in Construction 1 more carefully, as follows:
Construction 2.
Suppose the matrix $A$ is constructed as in Construction 1, where instead of $\mathbf{u}_{t}$ being an arbitrary unit vector (orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$), it is chosen specifically as the unit vector $\frac{1}{\|(I-N_{t}N_{t}^{\top})\boldsymbol{\xi}^{\prime}_{t}\|}(I-N_{t}N_{t}^{\top})\boldsymbol{\xi}^{\prime}_{t}$, where the columns of $N_{t}\in\mathbb{R}^{d\times q_{t}}$, $q_{t}\leq 2t-1$ are an orthogonal basis for $\mathbf{u}_{1},\ldots,\mathbf{v}_{t-1},\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$, and $\boldsymbol{\xi}^{\prime}_{t}\sim\mathcal{N}(\mathbf{0},I_{d})$ is an independent Gaussian.
By construction, $\mathbf{u}_{t}$ still satisfies the desideratum that it is a unit vector orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$. However, this particular construction gives us probabilistic control over the magnitude of its individual entries. Specifically, we have the following:
Lemma 5.3 (Good solution exists).
For $A$ as defined in Construction 2, it holds that
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}\geq\frac{\alpha}{\sqrt{d(T+1)}}~.$
Lemma 5.4 (Entries of $A$ are bounded).
Suppose $A$ is constructed as in Construction 2. Then there is a sufficiently large absolute constant $C>0$, such that for any $\delta\in(0,1)$ satisfying
$\frac{\log((n+d)/\delta)}{\sqrt{d}}\left(\alpha+\beta\sqrt{T}+1\right)\leq\frac{1}{C}~,$
it holds with probability at least $1-5nd\delta-\exp(-d/48)$ that $A\in[-1,+1]^{n\times d}$.
Note that the lemmas above are variants of Lemma 4.4 and Lemma 4.5 respectively. Combining them with Proposition 4.6 (which is unaffected by the additional constraint in Construction 2), and choosing $\alpha,\beta$ accordingly, the proof of Theorem 5.2 readily follows, as detailed in the appendix.
6 Discussion
In this paper, we studied the oracle complexity of solving matrix games based on the simplex. This well-studied problem class encapsulates tasks such as finding a linear separator and computing a Nash equilibrium in bilinear zero-sum games. By identifying distinct oracle models corresponding to either one-sided or two-sided multiplication by the matrix $A$, we were able to shed light on previous algorithmic approaches for this task, and to provide several new lower bounds.
For $\ell_{2}$/simplex games which correspond to the linear separability task, we first showed that a one-sided oracle leads to complexity no better than $\Omega(\gamma_{A}^{-2})$, attained by the classical Perceptron algorithm. Interestingly, this implies a separation between our oracle models, since accelerated methods utilizing two-sided oracle access can achieve a better rate of $\tilde{O}(\gamma_{A}^{-1})$. However, we showed that even with such improved methods, the achievable rate can be no better than $\tilde{\Omega}(\gamma_{A}^{-2/3})$. To the best of our knowledge, this is the first oracle complexity lower bound for this setting. Nonetheless, it still leaves a gap of $\tilde{O}(\gamma_{A}^{-1/3})$ between the upper and lower bounds. Closing this gap, either by designing better algorithms or by improving the lower bound, is an interesting open problem.
As for simplex/simplex and $\ell_{1}$/simplex games, we prove an $\tilde{\Omega}(\epsilon^{-2/3})$ iteration lower bound for computing an $\epsilon$-suboptimal solution, for any algorithm using a two-sided oracle, implying the same lower bound for computing an $\epsilon$-approximate Nash equilibrium. This is an exponential improvement over the previously known lower bound for this task due to Hadiji et al. (2024), leaving open a $\tilde{O}(\epsilon^{-1/3})$ gap compared to the best known upper bound of $\tilde{O}(\epsilon^{-1})$.
Another open problem is extending our lower bounds to randomized algorithms. Since our constructions required simulation of the algorithm’s responses, it is not immediately clear how this can be achieved. We note that previous extensions of oracle complexity lower bounds to randomized algorithms (Nemirovski and Yudin, 1983; Carmon et al., 2020a; Arjevani et al., 2023) crucially relied on non-linear modifications of the “hard” target function, which is not possible when we restrict ourselves to bilinear functions.
Acknowledgments
Appendix A Missing proofs from Section 4.2
A.1 Proof of Lemma 4.3
@@ Acknowledgments
Consider $\mathbf{w}=\frac{1}{\sqrt{T+1}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which is a unit vector since $\mathbf{u}_{1},\ldots,\mathbf{u}_{T+1}$ are orthonormal. We have
$A\mathbf{w}$~ $=~\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}+\mathbf{v}_{T}\mathbf{u}_{T+1}^{\top}\right)\left(\frac{1}{\sqrt{T+1}}\sum_{t=1}^{T+1}\mathbf{u}_{t}\right)$
@@
$=~\frac{1}{\sqrt{T+1}}\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})+\mathbf{v}_{T}\right)=~\frac{1}{\sqrt{T+1}}\cdot\mathbf{v}_{0}~=~\frac{\alpha}{\sqrt{T+1}}\cdot\mathbf{1}~.$
Therefore, $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}=\frac{\alpha}{\sqrt{T+1}}$ for this choice of unit vector $\mathbf{w}$, so the maximum over all unit vectors can therefore only be larger.
A.3 Proof of Lemma 4.5
Consider the vector $\mathbf{v}_{t}=\beta(I-M_{t}M_{t}^{\top})\boldsymbol{\xi}_{t}$ in the construction, for some $t$. Recalling that the matrix $M_{t}$ is a function of $\boldsymbol{\xi}_{1},\ldots,\mathbf{x}_{t-1}$, which are independent of $\boldsymbol{\xi}_{t}$, we have by Lemma C.1 that
$\Pr\left(\|\mathbf{v}_{t}\|_{\infty}\geq z~\middle|~\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{t-1}\right)~\leq~2n\exp\left(-\frac{z^{2}}{2\beta^{2}}\right)$
for all $z>0$. Since this holds for any realization of $\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{t-1}$, we can apply a union bound over all $t$ and get that
$\Pr\left(\max_{t\in[T]}\|\mathbf{v}_{t}\|_{\infty}\geq z\right)~\leq~2Tn\exp\left(-\frac{z^{2}}{2\beta^{2}}\right)~.$
Choosing $z$ appropriately, this can equivalently be phrased as follows: For any $\delta\in(0,1)$, with probability at least $1-\delta$, $\max_{t\in[T],l\in[n]}|v_{t,l}|\leq\beta\sqrt{2\log(2Tn/\delta)}$. Under this event, and recalling that $v_{0,l}=\alpha$ by construction, it holds for any $l\in[n]$ that
$\|A_{l}\|^{2}$~ $=~\left\|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})\mathbf{u}_{j}^{\top}+v_{T,l}\mathbf{u}_{T+1}^{\top}\right\|^{2}$
$\stackrel{{\scriptstyle(1)}}{{=}}~\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})^{2}+v_{T,l}^{2}~\stackrel{{\scriptstyle(2)}}{{=}}~\sum_{j=1}^{T}2(v_{j-1,l}^{2}+v_{j,l}^{2})+v_{T,l}^{2}$
$\leq~2v_{0,l}^{2}+4\sum_{j=1}^{T}v_{j,l}^{2}~\leq~2\alpha^{2}+4T\beta^{2}\cdot 2\log(2Tn/\delta)~,$
where $(1)$ is because $\mathbf{u}_{1},\ldots,\mathbf{u}_{T+1}$ are orthogonal unit vectors, and $(2)$ is because $(x-y)^{2}\leq 2x^{2}+2y^{2}$ for any $x,y\in\mathbb{R}$. By the assumption stated in the lemma, this expression is at most $1$ as required.
A.4 Proof of Proposition 4.6

### v4

@@ 4 Warm up: Linear separability with a one-sided oracle
6 Applications
In this section, we discuss some key implications of our results.
6.1 Nash equilibria
A prominent problem class of interest is that of simplex/simplex games, which are well-known to model zero-sum games. Given a matrix with bounded entries $A\in[-1,1]^{n\times d}$, or equivalently $\|A\|_{1\to\infty}\leq 1$, the problem is to solve
$\max_{\mathbf{w}\in\Delta^{d-1}}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~.$ (5)
Recall that for this problem, a two-sided oracle is exactly equivalent to a first-order (or gradient) oracle, which returns the gradient of $(\mathbf{p},\mathbf{w})\mapsto\mathbf{p}^{\top}A\mathbf{w}$. Thus, a lower bound under the two-sided oracle model is precisely a lower bound for solving zero-sum games using a first-order oracle.
To apply Theorem 5.1, which holds for $\ell_{p}$ balls, we provide a reduction from simplex/simplex games to $\ell_{1}$/simplex games, showing that any lower bound for the latter implies a lower bound for the former. In other words, any algorithm for solving Problem (5) (over the simplex) can be readily converted to an algorithm for solving $\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$ (in a slightly modified dimension).
@@ 6.1 Nash equilibria
The formal proof is deferred to the proof section. In a nutshell, the idea is that given a matrix $A\in[-1,+1]^{n\times d}$ and an algorithm for a simplex/simplex matrix game, we can feed the algorithm with the matrix $(A;-A)\in\mathbb{R}^{n\times 2d}$, and convert the algorithm’s output $\mathbf{w}$ (a vector in the (2d1)(2d-1)-simplex) to a $d$-dimensional vector $\mathbf{w}^{\prime}$ in the $\ell_{1}$ unit ball, so that $A\mathbf{w}^{\prime}=(A;-A)\mathbf{w}$, leading to similar guarantees. We note that the reduction does require us to modify the matrix width $d$ by a factor of $2$, but this will not affect our lower bound by more than a small constant factor (as the bound we will show applies to any sufficiently large $d$). We also note that the computational complexity of the two algorithms in the reduction are essentially identical.
Combining Theorem 5.1 for $p=1$, with Proposition 6.1, results in the following corollary:
Corollary 6.2.
No deterministic first-order algorithm can guarantee returning a $\epsilon$-approximate Nash equilibrium of a matrix $A\in[-1,1]^{n\times d}$ (for sufficiently large $d,n\in\mathrm{poly}(1/\epsilon)$), using less than $\tilde{\Omega}(\epsilon^{-2/3})$ calls to a two-sided oracle $\mathcal{O}^{A}_{2}$. Moreover, as this lower bound is nearly-matched by Karmarkar et al. (2026), our result settles the first-order oracle complexity of this task (up to log factors).
Our lower bound for approximating Nash equilibria is nearly-matched by Karmarkar et al. (2026), hence our result settles the oracle complexity of this task (up to log factors) in the natural model of matrix-vector multiplications.
6.2 Linear separability
As previously mentioned, $\ell_{2}$/simplex games model the task of finding a linear separator: Given a linearly separable classification dataset $(\mathbf{x}_{l},y_{l})_{l=1}^{n}\subset\mathbb{R}^{d}\times\{\pm 1\}$ with bounded norm $\max_{l\in[n]}\|\mathbf{x}_{l}\|_{2}\leq 1$, consider the data matrix $A\in\mathbb{R}^{n\times d}$ with $A_{l}:=y_{l}\mathbf{x}_{l}$. Then $\|A\|_{2\to\infty}\leq 1$, and the problem
$\max_{\mathbf{w}\in\mathbb{R}^{d}:\|\mathbf{w}\|_{2}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$
corresponds to finding a linear separator with maximal margin $\gamma_{A}$.
To the best of our knowledge, no oracle complexity lower bounds for this task that appear in the current literature apply to accelerated algorithms, which as we showed in Theorem 4.1, provably require two-sided multiplications of the data matrix $A$. Applying Theorem 5.1 for $p=2$ results in the following corollary:
Corollary 6.3.
No deterministic algorithm can guarantee returning a linear separator of a classification dataset as above (for sufficiently large $d,n\in\mathrm{poly}(1/\gamma_{A})$), using less than $\tilde{\Omega}(\gamma_{A}^{-2/3})$ calls to a two-sided oracle $\mathcal{O}^{A}_{2}$ of the data matrix.
Our lower bound for finding a linear separator is nearly-matched by Karmarkar et al. (2026), hence our result settles the oracle complexity of this task (up to log factors) in the natural model of matrix-vector multiplications.
6.3 Combinatorial applications
Another problem class which received notable interest in the literature is that of $\ell_{\infty}$/simplex games, also referred to as “box-simplex games”, of the form
$\max_{\mathbf{w}\in[-1,1]^{d}}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~.$
Solving such games appears as an important subroutine in a variety of applications, mostly of combinatorial nature, such as the computing a graph’s max-flow (Sherman, 2017), bipartite matching (Assadi et al., 2022), densest subgraph (Boob et al., 2019; Nguyen and Ene, 2024), as well as other computational tasks such as optimal transport (Jambulapati et al., 2019).
Sherman (2017) showed that given a $\ell_{\infty}$/simplex game with matrix $A$ satisfying $\|A\|_{\infty\to\infty}\leq 1$, an $\epsilon$-suboptimal solution can be returned using $O(\log(d)\epsilon^{-1}\log(\epsilon^{-1}))$ two-sided queries. To the best of our knowledge, no oracle complexity lower bound is known for this task. Applying Theorem 5.1 for $p=\infty$ results in the following corollary:
Corollary 6.4.
No deterministic algorithm can guarantee returning an $\epsilon$-suboptimal solution to a $\ell_{\infty}$/simplex game with respect to a $n\times d$ matrix $A$ (for sufficiently large $d,n\in\mathrm{poly}(1/\epsilon)$), using less than $\tilde{\Omega}(\epsilon^{-2/3})$ calls to a two-sided oracle $\mathcal{O}^{A}_{2}$.
Notably, as opposed to the previously discussed applications for $p\in\{1,2\}$, this result leaves open a $\tilde{O}(\epsilon^{-1/3})$ gap between the best-known upper bound and the lower bound we have established, leaving open an interesting problem for future work.
7 Proofs
7.1 Proof of Lemma 4.3
We start by noting that the case $p=2$ simply follows from orthonormality of $A$’s rows, since in that case it holds deterministically that
$\frac{\|A_{l}\|_{2}^{2}}{\|\sum_{t=1}^{T}A_{t}\|_{2}}=\frac{\|A_{l}\|_{2}^{2}}{\sqrt{\sum_{t=1}^{T+1}\|A_{t}\|_{2}^{2}}}=\frac{1}{\sqrt{T+1}}~.$
We move on to prove the general probabilistic case, and denote $\mathbf{g}_{t}:=\Pi_{t}(\boldsymbol{\xi}_{t})$ so that $A_{t}=\frac{\mathbf{g}_{t}}{\|\mathbf{g}_{t}\|_{q}}$. Note that projecting a Gaussian results in a Gaussian with smaller (or equal) variance, and that in this case $\boldsymbol{\xi}_{t}$ is projected onto a subspace of dimension at least $d-2t\geq d-2T\geq\frac{1}{2}d$. We can apply a standard Gaussian concentration bound, concretely the Laurent–Massart bound, to ensure that with probability at least $\frac{3}{4T}:$
$\|\mathbf{g}_{t}\|_{2}^{2}\geq\frac{1}{2}d-2\sqrt{d\log(4T)/2}\geq\frac{1}{4}d~,$
an event which further implies that
$\|\mathbf{g}_{t}\|_{q}\geq d^{1/q-1/2}\|\mathbf{g}_{t}\|_{2}\geq\frac{1}{2}d^{1/q}~.$
Furthermore, a standard Gaussian tail bound $\Pr(|g_{t,i}|>z)\leq 2\exp(-z^{2}/2)$ implies via a union bound that with probability at least $\frac{3}{4T}:$
$\|\mathbf{g}_{t}\|_{q}\leq d^{1/q}\sqrt{2\log(16Td)}~~~\text{and}~~~\|\mathbf{g}_{t}\|^{2}_{2}\leq 2d\log(8Td)~.$
Union bounding over $t$, we get that with probability at least $\frac{1}{2}:$
$\|A_{l}\|_{2}^{2}=\frac{\|\mathbf{g}_{l}\|_{2}^{2}}{\|\mathbf{g}_{l}\|_{q}^{2}}\geq\frac{\frac{1}{4}d}{2d^{2/q}\log(16Td)}=\frac{d^{1-2/q}}{8\log(16Td)}~,$
and also by using the orthogonality of $A$’s rows,
$\left\|\sum_{t=1}^{T+1}A_{t}\right\|_{p}$ $\leq d^{1/p-1/2}\left\|\sum_{t=1}^{T+1}A_{t}\right\|_{2}=d^{1/p-1/2}\sqrt{\sum_{t=1}^{T+1}\|A_{t}\|_{2}^{2}}=d^{1/p-1/2}\sqrt{\sum_{t=1}^{T+1}\frac{\|\mathbf{g}_{t}\|_{2}^{2}}{\|\mathbf{g}_{t}\|_{q}^{2}}}$
$\leq d^{1/p-1/2}\sqrt{(T+1)\cdot\frac{2d\log(8Td)}{\frac{1}{4}d^{2/q}}}$
$\leq 4d^{1/p-1/q}\sqrt{(T+1)\log(8Td)}~.$
Thus overall,
$\frac{\|A_{l}\|_{2}^{2}}{\left\|\sum_{t=1}^{T+1}A_{t}\right\|_{p}}$ $\geq\frac{d^{1-2/q}}{8\log(16Td)\cdot 4d^{1/p-1/q}\sqrt{(T+1)\log(8Td)}}$
$=\frac{1}{32\sqrt{T+1}\cdot\log(16Td)\sqrt{\log(8Td)}}~,$
where we used the fact that $1-\frac{2}{q}-\frac{1}{p}+\frac{1}{q}=1-\frac{1}{q}-\frac{1}{p}=0$. Crudely bounding $\log(8Td)\leq\log(16Td)\leq\log(16d^{2})=2\log(4d)$ completes the proof.
7.2 Proof of Lemma 5.3
@@ 7 Proofs
Consider $\mathbf{w}=\frac{1}{\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{p}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which by construction satisfies $\|\mathbf{w}\|_{p}\leq 1$. Since $\mathbf{u}_{1},\ldots,\mathbf{u}_{T+1}$ are orthonormal, we have
$A\mathbf{w}$ $=\frac{1}{\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{p}}\cdot\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}+\mathbf{v}_{T}\mathbf{u}_{T+1}^{\top}\right)\left(\sum_{t=1}^{T+1}\mathbf{u}_{t}\right)$
$=\frac{1}{\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{p}}\cdot\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})+\mathbf{v}_{T}\right)$
$=\frac{1}{\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{p}}\cdot\mathbf{v}_{0}$
$=\frac{\alpha}{\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{p}}\cdot\mathbf{1}~.$
Thus, each entry of $A\mathbf{w}$ equals $\alpha/\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{p}$, and since this holds for the $\mathbf{w}$ we have chosen, the maximum of $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}$ over all vectors $\mathbf{w}$ in the unit $\ell_{p}$ ball can only be larger. Therefore, it remains to upper bound the quantity $\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{p}$.
For $p\in[1,2]$, it suffices to recall that $\mathbf{u}_{1},\dots,\mathbf{u}_{T+1}$ are orthonormal, and by comparing the $\ell_{p}$ and $\ell_{2}$ norms we see that
$\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{p}\leq d^{\frac{1}{p}-\frac{1}{2}}\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{2}=d^{\frac{1}{p}-\frac{1}{2}}\sqrt{T+1}\leq 2d^{\frac{1}{p}}\sqrt{T}~,$
proving the first lower bound in the lemma (deterministically).
For $p\in(2,\infty)$ we use a probabilistic argument. Note that for any index $i\in[d]$, the sequence $u_{1,i},\dots,u_{T+1,i}$ is constructed so that for any $t\in[T+1]:~u_{t,i}\mid u_{1,i},\dots,u_{t-1,i}$ is a mean-zero sub-Gaussian random variable with sub-Gaussian norm bounded by $O(1/d)$ (note that the projected dimension is still at least on the order of $d$ by assumption that $T\ll d$). Hence, a martingale bound (cf. Shamir, 2011) ensures that with probability at least $1-\delta/d$ it holds that $|\sum_{t\in[T+1]}u_{t,i}|\leq C\sqrt{T\log(d/\delta)/d}$ for some universal constant $C>0$. Union bounding over $i\in[d]$ so that the bound holds for all $i$ with probability at least $1-\delta$, under this probable event we get
$\left\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\right\|_{p}$ $=\left(\sum_{i\in[d]}\left|\sum_{t\in[T+1]}u_{t,i}\right|^{p}\right)^{1/p}\leq\left(\sum_{i\in[d]}C^{p}\left(\frac{T\log(d/\delta)}{d}\right)^{p/2}\right)^{1/p}$
$=Cd^{1/p-1/2}\sqrt{T\log(d/\delta)}~,$
proving the claimed bound for $p\in(2,\infty)$.
For $p=\infty$, under the same probable event as before, it holds that $\|\sum_{t=1}^{T+1}\mathbf{u}_{t}\|_{\infty}=\max_{i\in[d]}|\sum_{t\in[T+1]}u_{t,i}|\leq C\sqrt{T\log(d/\delta)/d}$ as well.
7.4 Proof of Lemma 5.5
Recall that $\|A\|_{p\to\infty}=\max_{l\in[n]}\|A_{l}\|_{q}$ where $q\in[1,\infty]$ is the dual exponent satisfying $1/p+1/q=1$, and so we set out to bound the $q$-norm of $A$’s rows. For any $l\in[n]$, it holds that
$\|A_{l}\|_{q}$ $=\left\|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})\mathbf{u}_{j}^{\top}+v_{T,l}\mathbf{u}_{T+1}^{\top}\right\|_{q}$
$\leq\left\|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})\mathbf{u}_{j}^{\top}\right\|_{q}+\|v_{T,l}\mathbf{u}_{T+1}^{\top}\|_{q}~.$ (6)
We turn to bound the terms above. For $l\in[n],i\in[d]$, we denote $D^{j}_{l}:=v_{j-1,l}-v_{j,l}$ and $S_{l,i}:=\sum_{j=1}^{T}D^{j}_{l}u_{j,i}$. Recall (as in the proof of Lemma 5.4) that each $u_{j,i}$ is a sub-Gaussian, mean-zero random variable with sub-Gaussian norm bounded by $O(1/d)$, and therefore conditioned on $\mathbf{v}_{0},\dots,\mathbf{v}_{T}$, the random variable $S_{l,i}$ has sub-Gaussian norm at most $O(\frac{1}{d})\cdot\sum_{j=1}^{T}(D_{l}^{j})^{2}$ (this follows from a basic summation property for sub-Gaussian norms, cf. Vershynin, 2018). Therefore, there exists a universal constant $C_{1}>0$ such that for any $z>0:$
The numerical inequality $(a-b)^{2}\leq 2(a^{2}+b^{2})$ implies that $\sum_{j=1}^{T}(D^{j}_{l})^{2}\leq 4\sum_{j=0}^{T}v_{j,l}^{2}=4\alpha^{2}+4\sum_{j=1}^{T}v_{j,l}^{2}$, and further note that while $v_{1,l},\dots,v_{T,l}$ are not independent, it does hold that conditionally $v_{t,l}\mid v_{0,l},\dots v_{t-1,l}\sim\mathcal{N}(0,\sigma_{t,l}^{2})$ for some $\sigma_{t,l}\leq\beta$ since projecting a Gaussian cannot increase its variance. Therefore, Lemma 7.5 ensures that with probability at least $1-\delta:~\sum_{j=1}^{T}v_{j,l}^{2}\leq 8\beta^{2}(T+\log(1/\delta))$, so via a union bound, we get that with probability at least $1-2\delta$ it holds for all $z>0$ that
Equivalently, choosing $z$ by equating the above to $\delta$, slightly simplifying the expression and union bounding over $l\in[n]$ and $i\in[d]$, we see that for some sufficiently large absolute constant $C>0$ it holds with probability at least $1-\delta$ that
@@
$|S_{l,i}|\leq C\sqrt{\frac{\log(nd/\delta)}{d}}\left[\alpha+\beta(\sqrt{T}+\sqrt{\log(nd/\delta)})\right]~,\text{~~for all }l\in[n],\,i\in[d]~.$ (7)
Under this event, for any $q<\infty$, it holds that
$\left\|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})\mathbf{u}_{j}^{\top}\right\|_{q}$ $=\left(\sum_{i=1}^{d}\left|S_{l,i}\right|^{q}\right)^{1/q}$ (8)
$\leq\left(\sum_{i=1}^{d}\left(C\sqrt{\frac{\log(nd/\delta)}{d}}\left[\alpha+\beta(\sqrt{T}+\sqrt{\log(nd/\delta)})\right]\right)^{q}\right)^{1/q}$
$=Cd^{1/q-1/2}\sqrt{\log(nd/\delta)}\left[\alpha+\beta(\sqrt{T}+\sqrt{\log(nd/\delta)})\right]~,$
thus bounding the first term in Eq. (6). We similarly derive the bound for $q=\infty$ by simply applying Eq. (7), since $\|\sum_{j=1}^{T}(v_{j-1,l}-v_{j,l})\mathbf{u}_{j}^{\top}\|_{\infty}=\max_{i\in[d]}|S_{l,i}|$.
To bound the second term in Eq. (6), we apply Lemma 7.4 twice to get that with probability at least $1-2\delta-\exp(-d/48):$
$\|\mathbf{v}_{T}\|_{\infty}~\leq~\beta\sqrt{2\log(2n/\delta)}\text{~~~~~and~~~~~}\|\mathbf{u}_{T+1}\|_{\infty}\leq\sqrt{8\log(2d/\delta)/d}~,$
and under this event, for any $l\in[n]$ it holds that
$\|v_{T,l}\mathbf{u}_{T+1}^{\top}\|_{q}$ $=|v_{T,l}|\cdot\|\mathbf{u}_{T+1}^{\top}\|_{q}\leq\|\mathbf{v}_{T}\|_{\infty}\cdot d^{1/q}\|\mathbf{u}_{T+1}^{\top}\|_{\infty}$
$\leq 4\beta d^{1/q-1/2}\sqrt{\log(2n/\delta)\log(2d/\delta)}~.$ (9)
Overall, plugging Eq. (8) and Eq. (9) into Eq. (6) and noting that the second summand is dominated by the first (for sufficiently large constant $C>0$), rescaling $\delta$ by a factor of $3$ and slightly simplifying completes the proof.
7.5 Proof of Proposition 5.6
We start by introducing notation that will be used throughout the proof. Let $M_{t}\in\mathbb{R}^{n\times\mathrm{dim}(\mathrm{span}(\mathbf{v}_{1},\dots,\mathbf{v}_{t-1},\mathbf{p}_{1},\dots,\mathbf{p}_{t}))}$ be a matrix whose columns are an orthonormal basis for $\mathrm{span}(\mathbf{v}_{1},\dots,\mathbf{v}_{t-1},\mathbf{p}_{1},\dots,\mathbf{p}_{t})$, and note that $\mathbf{v}_{t}=\beta(I-M_{t}M_{t}^{T})\boldsymbol{\xi}_{t}$.